An amount of $1,000 is deposited into a bank account that pays 4% interest compounded once a year. If there are no other withdrawals or deposits, what will be the balance of the account after 3 years?

Answer by mananth(16949)


Find Amount
Principal P = 1000
Amount= A
years=t 3.00
compounded 1 times a year n
Rate = 4.00 0.04 0.04
Amount = P*((n+r)/n)^n*t

Amount = = 1000 *( 1 + 0.04 )^ 1 * 3.00
Amount = 1000 *( 1 + 0.04 )^ 3
1000 *( 1.04 )^ 3
Amount = 1124.86